OpenStudy (anonymous):

what is 523% of 952.

OpenStudy (anonymous):

523% of 952 = (523/100)*952 = 4978.96

OpenStudy (anonymous):

First, write 523% as the fraction 523/100. Now we solve the same way we would if they asked us what is 1/2 of x - by multiplying.

hero (hero):

Hint: % = "per 100" of = multiply So: \[\frac{523}{100} \times 952\]

OpenStudy (anonymous):

thats what i have also

OpenStudy (anonymous):

thank you both